# Learning Hidden Physics and System Parameters with Deep Operator Networks

### v1

@@ 4          Branch network: process sensor inputs $u(x^{s},t^{s})$
5          Trunk network: process coordnates $(x,t)$
@@
7       for each batch of sensor data do
@@ 10             Compute loss
11             $\mathcal{L}(\theta)=\frac{1}{N}\sum_{i=1}^{N}||\text{prediction}_{i}-u_{i}(x^{s},t^{s})||^{2}$
@@ 27             Compute combine loss
28             $\mathcal{L}^{\text{pde}}=\frac{1}{N}\sum_{i=1}^{N}||\mathcal{N}(\mathcal{X},u_{\text{pred}_{i}},\xi_{\text{pred}_{i}})-g_{i}||^{2}$   physics loss
29             $\mathcal{L}^{\text{data}}=\frac{1}{N}\sum_{i=1}^{N}||u_{\text{pred}_{i}}-u(x^{s},t^{s})||^{2}$   data loss
@@ 31             Back-propagate and update $\bm{\theta}$ and $\bm{\phi}$ using Adam
We have consider the Loss tolerance $\epsilon$ of $\mathcal{O}(10^{-3})$. To measure the deviation of the predicted solution from reference we use a metric known as relative $L_{2}$ Error which is defined as,
$\text{relative}\ L_{2}\ \text{Error}=\frac{\sqrt{\sum_{i=1}^{N}|u_{i}-u_{i}^{*}|^{2}}}{\sqrt{\sum_{i=1}^{N}|u_{i}^{*}|^{2}}}$ (7)
where $u^{*},\;u$ are the reference and the predicted solution respectively. We also define, the error distribution by $\mu\pm\sigma$, where $\mu,\sigma$ are mean and standard deviation of rel. $L_{2}$ errors.
@@ 3 Illustrative Problems
To examine our approach, we perform tests on two examples including the Reaction Diffusion equation and Burger’s equation. The input function $f$ acts as the source term for the reaction-diffusion system and initial condition for the burgers’ equation.
@@ 3.1 Reaction Diffusion System
We first examine our approaches using a reaction-diffusion equation, which models various physical phenomena in nature such as oscillating chemical concentrations and spatial waves as seen in the Belousov-Zhabotinsky reaction, population dynamics, pattern formation in biological systems and morphology in crystals and alloys [44]. The following equation describes the spatiotemporal evolution of the system and is given by:
$\begin{split}&\frac{\partial u}{\partial t}(x,t)=D\frac{\partial^{2}u}{\partial x^{2}}(x,t)+Ku^{2}(x,t)+f(x)\text{ on }\ \Omega:(x,t)\in[0,1]^{2},\\ &IC:u(x,0)=0,\\ &BC:u(0,t)=u(1,t)=0,\\ \end{split}$ (8)
@@
    with RBF kernel $k_{l}(x_{1},x_{2})=exp(-||x_{1}-x_{2}||^{2}/(2l^{2}))$ with a length-scale parameter $l>0$. The $l$ determines the oscillatory nature of the sampled function, a lower value of $l$ leads to higher oscillations. We consider GRFs with RBF kernel length-scale of 0.2, unlike sine basis functions GRFs are not zero at boundaries.
@@
To train the framework, the collocation points in the differential equation, initial conditions, and boundary conditions are defined as $N_{Coll}$, $N_{ic}$, and $N_{bc}$, respectively. These points are randomly generated for each sample using Latin hypercube sampling [45], creating synthetic domain data that can vary at each gradient descent step. The training process is shown in Algorithm 1. For this problem, we use $N_{Coll}=2000$, $N_{ic}=200$, and $N_{bc}=250$. To determine the optimal amount of known labeled dataset required, we experiment with data point sets $N_{d}=\{200,500\}$ and training sample sets $N_{train}=\{50,100,200,500\}$, with a fixed number of test samples, $N_{test}=1000$.
The branch net takes the source term discretized at 101 spatial sensor locations, while the trunk net takes the spatial and temporal locations while the hidden physics MLP takes the solution $u$ and the corresponding spatial gradients, $u,u_{x},$ and $u_{xx}$ as inputs to output the prediction $\mathcal{N}$ for RHS in Equation 9. The optimal model was found using training size of $N_{train}=500,N_{d}=500$. The loss curve over the training process is shown in Figure 1(a). The mean test relative $L_{2}$ error for each of the cases is presented in Table 1.
@@
Sine basis $0.01742\pm 0.007142$
GRF $0.14039\pm 0.09857$
Modified GRF $0.04028\pm 0.03121$
Using sine basis functions for the source term yields the highest prediction accuracy among tested methods. Figure 3(a) illustrates the mean test error over various sample sizes and training data points, while Figure 3(b) shows the distribution of test errors for this optimal model. We further evaluate the model’s predictive capability on unseen sine basis source terms as defined in Equation 10. As shown in Figure 5, the model accurately predicts both the solution $u$ and the hidden physics terms $\mathcal{N}$, aligning closely with ground truth values. The inferred hidden physics term is given by
@@
Additionally, we assess the model’s performance on GRFs by varying the length-scale parameter $l$ of the RBF kernel with values $l\in\{0.1,0.15,0.20,0.40,0.60\}$. As expected, Table 2 demonstrates that prediction error increases as the length scale decreases. However, for out-of-distribution functions with boundary values approaching zero, the model maintains relatively higher accuracy compared to cases with non-zero boundary conditions.
Figure 3: Performance analysis for Reaction diffusion equation: (a) Mean test error over varying $N_{\text{train}}$. (b) Test error distribution of optimal model that achieved an error $0.01742\pm 0.007142$ with $N_{train}=500$, $N_{d}=500$.
@@
0.10 0.1953 $\pm$ 0.1263
0.15 0.09055 $\pm$ 0.07274
0.20 0.04028 $\pm$ 0.03121
0.40 0.01720 $\pm$ 0.01026
0.60 0.01593 $\pm$ 0.01077
Figure 4: Performance analysis for GRFs and Modified GRFs input function spaces: (a) GRFs basis $l=0.2$, Test error = 0.14039 $\pm$ 0.09857. (b) Modified GRFs basis $l=0.2$, Test error = 0.04028 $\pm$ 0.03121
(a) Sample 1: solution field accuracy comparison, relative $L_{2}$ error = 0.01639.
(b) Sample 1: hidden physics solution comparison, relative $L_{2}$ error = 0.03949.
(c) Sample 2: solution field accuracy comparison, relative $L_{2}$ error = 0.01960.
(d) Sample 2: hidden physics solution comparison, relative $L_{2}$ error = 0.02435.
Figure 5: Reaction diffusion equation: Comparison of reference and predicted results for two representative test samples with input function from sine function space.
(a) Sample 1: solution field accuracy comparison, relative $L_{2}$ error = 0.02733.
(b) Sample 1: hidden physics solution comparison, relative $L_{2}$ error = 0.34308.
(c) Sample 2: solution field accuracy comparison, relative $L_{2}$ error = 0.0322.
(d) Sample 2: hidden physics solution comparison, relative $L_{2}$ error = 0.23023.
Figure 6: Reaction diffusion equation: Comparison of reference and predicted results for two representative test samples with input function from GRF function space $l=0.2$ (nonzero at boundaries)
(a) Sample 1: solution field accuracy comparison, relative $L_{2}$ error = 0.01193.
(b) Sample 1: hidden physics solution comparison, relative $L_{2}$ error = 0.05668.
(c) Sample 2: solution field accuracy comparison, relative $L_{2}$ error = 0.02303.
(d) Sample 2: hidden physics solution comparison, relative $L_{2}$ error = 0.07587.
Figure 7: Reaction diffusion equation: Comparison of reference and predicted results for two representative test samples with input function from Modified GRF function space $l=0.2$ (zero at boundaries)
@@
The neural operator framework introduced in Section 2.2 is employed to characterize the system, specifically to identify the diffusion coefficient of the governing PDE using solution values at selected spatial and temporal locations.
@@
(a) Sample 1: Relative $L_{2}$ error of $u(x,t)$ = $0.027$. For this case, $\nu_{\text{true}}=0.027$ while $\nu_{\text{predicted}}=0.027$.
(b) Sample 2: Relative $L_{2}$ error of $u(x,t)$ = $0.029$. For this case, $\nu_{\text{true}}=0.036$ while $\nu_{\text{predicted}}=0.034$.
@@
Figure 9: Reaction Diffusion Equation: Distribution of absolute error of $D$ when compared to the $D$ used to generate the ground truth using the finite difference solver. The framework approximates $D$ with a error of $8\times 10^{-3}\pm 0.008$.
3.2 Burgers equation
@@
$\begin{split}&\frac{\partial u}{\partial t}(x,t)=\nu\frac{\partial^{2}u}{\partial x^{2}}(x,t)-u\frac{\partial u}{\partial x}(x,t)\text{ on }\ \Omega:(x,t)\in[0,1]^{2},\\ \text{IC: }&u(x,0)=f(x),\\ \textnormal{BC: }&u(0,t)=u(1,t)\text{ and }\frac{\partial u}{\partial x}(0,t)=\frac{\partial u}{\partial x}(1,t),\\ \end{split}$ (13)
@@
where the $\mathcal{N}$ denotes the hidden physics that is a function of the solution field $u(x,t)$ and its spatial derivatives. In this problem, our aim is to learn the solution operator, $\mathcal{G}_{\theta}$ that maps the initial condition, $f(x,0)$ to the solution field, $u(x,t)$ using the partial known physics as well as sparse labeled dataset. In this case, we have considered fixed viscosity for the system, $\nu=0.01$. We solved equation (13) using a finite difference solver on a uniform grid ($\delta x=\delta t=0.01$), resulting in a $101\times 101$ grid spanning the $x$-$t$ domain with initial conditions sampled from a Gaussian Random Field (GRF).
The experiments were conducted with varying numbers of training samples, $N_{train}\in\{50,100,200,500,1000\}$ and the size of labeled dataset, $N_{d}\in\{200,500\}$, where $N_{d}$ denotes the number of locations per sample where the value of solution field is known. For testing, we used a fixed set of $N_{test}=1,000$ samples. Figure 10(a) illustrates the mean test error across different combinations of sample sizes and training data points. The optimal performance was achieved with $N_{train}=1000$ and $N_{d}=500$, resulting in a mean test relative $L_{2}$ error of 0.06140. The loss curve over the training process is shown in Figure 1(b).
The distribution of test errors for this optimal configuration is presented in Figure 10(b). As demonstrated in Figure 11, the model accurately predicts both the solution $u$ and the hidden physics terms when compared to ground truth values. Therefore, the model successfully approximated the unknown hidden term as
@@
Figure 10: Performance analysis for Burger’s equation: (a) Mean test error over varying $N_{\text{train}}$. (b) Test error distribution of optimal model that achieved an error $0.06140\pm 0.03510$ with $N_{train}=1000$, $N_{d}=500$.
(a) Sample 1: solution field accuracy comparison, relative $L_{2}$ error = 0.02134.
(b) Sample 1: hidden physics solution comparison, relative $L_{2}$ error = 0.118278.
(c) Sample 2: solution field accuracy comparison, relative $L_{2}$ error = 0.03909.
(d) Sample 2: hidden physics solution comparison, relative $L_{2}$ error = 0.15396.
Figure 11: Burgers’ equation: Comparison of reference and predicted results for two representative test samples.
@@
We are now interested in employing the neural operator framework discussed in Section 2.2 to characterize the system, more specifically to identify the viscosity of the model while being provided with the solution values at a few spatial and temporal locations.

### v2

@@ 4     Branch network: process sensor inputs $u(x^{s},t^{s})$
5     Trunk network: process coordinates $(x,t)$
@@
7 for for each batch of K system realizations do
@@ 10     Compute loss
11     $\mathcal{L}(\theta)=\frac{1}{K}\sum_{i=1}^{K}\left\lVert\text{prediction}_{i}-u_{i}(x^{s},t^{s})\right\rVert^{2}$
@@ 27    Compute combine loss
[…]
@@ 31      Back-propagate and update $\bm{\theta}$ and $\bm{\phi}$ using Adam
Algorithm 2 presents our two-stage training procedure for learning the parameter identification operator. The algorithm operates on a training dataset comprising multiple system realizations, where each realization corresponds to distinct parameter values and yields sensor measurements at fixed spatial locations over time. During training, the networks process mini-batches of these realizations, optimizing the loss functions through standard stochastic gradient descent until convergence.
@@ 3 Illustrative Problems
To examine our approach, we perform tests on four sets of examples including the Reaction Diffusion equation, Burger’s equation, Heat equation and Helmholtz Equation. The network architectures were systematically selected following established DeepONet guidelines Lu et al. [2021], with branch and trunk networks employing 3-4 hidden layers of 64-128 neurons. Hyperparameter sweep analysis demonstrates robust performance across reasonable architectural variations. Complete details are provided in Appendix A (Table A.1). For each problem, we demonstrate the applicability of our frameworks for hidden physics discovery, parameter identification, or both depending on the problem configuration.
@@ 3.1 Reaction Diffusion System
We first examine our approaches using a reaction-diffusion equation, which models various physical phenomena in nature such as oscillating chemical concentrations and spatial waves as seen in the Belousov-Zhabotinsky reaction, population dynamics, pattern formation in biological systems and morphology in crystals and alloys Smoller [2012]. The following equation describes the spatiotemporal evolution of the system and is given by:
[…]
@@
[…]
@@
To train the framework, the collocation points in the differential equation, initial conditions, and boundary conditions are defined as $N_{Coll}$, $N_{ic}$, and $N_{bc}$, respectively. These points are randomly generated for each sample using Latin hypercube sampling McKay et al. [1979], creating synthetic domain data that can vary at each gradient descent step. The training process is shown in Algorithm 1. For this problem, we use $N_{Coll}=2000$, $N_{ic}=200$, and $N_{bc}=250$. To determine the optimal amount of known labeled dataset required, we experiment with data point sets $N_{d}=\{200,500\}$ and training sample sets $N_{train}=\{50,100,200,500\}$, with a fixed number of test samples, $N_{test}=1000$.
The branch net takes the source term discretized at 101 spatial sensor locations, while the trunk net takes the spatial and temporal locations while the hidden physics MLP acts as a surrogate model for the right-hand side (RHS) of Equation 10. It accomplishes this by taking the solution $u$ and its spatial gradients, $u,u_{x},u_{xx}]$ as inputs and producing the prediction $\mathcal{N}$. The optimal model was found using training size of $N_{train}=500,N_{d}=500$. The loss curve over the training process is shown in Figure 1(a). The test mean absolute error for each of the cases is presented in Table 1, The solution is predicted within training time domain unless specifically mentioned.
@@
Sine basis $0.00871\pm 0.00424$
GRF $0.03918\pm 0.02095$
Modified GRF $0.01653\pm 0.00856$
Using sine basis functions for the source term yields the highest prediction accuracy among tested methods. Figure 4(a) illustrates the mean test error over various sample sizes and training data points, while Figure 4(b) shows the distribution of test errors for this optimal model with sine basis input function. We further evaluate the model’s predictive capability by using the learned hidden physics MLP $\mathcal{N}(x,t_{n};f)$ to solve the governing PDE via time integration beyond the training data (temporal extrapolation). Starting from the initial condition $u_{0}$, we apply forward Euler integration with the discovered physics term. For the Reaction-Diffusion system, the update rule is $u_{n+1}(x)=u_{n}(x)+(\mathcal{N}(x,t_{n};f)+f(x))\cdot dt$, while for Burgers’ equation it simplifies to $u_{n+1}(x)=u_{n}(x)+\mathcal{N}(x,t_{n};f)\cdot dt$. We compare two solution approaches: the direct DHPO prediction $u_{\text{pred}}$ (querying the trained DeepONet at future times) versus the time-integrated solution $u_{\text{hpm}}$ (obtained by stepping forward using the learned physics). We assess both approaches on unseen sine basis source terms as defined in Equation 11.
As shown in Figure 3, the model accurately predicts the DHPO solution ($u_{pred}$) and DHPO MLP integrated solution ($u_{hpm}$) at the training region. We also observe, in extrapolation region, the model accurately predicts the solution up to a specific time threshold, the accuracy in integrated solution is higher than DHPO solution, Hence we can infer that the Hidden Physics MLP is accurately capturing the dynamics. Figure 6 shows the contours of reference and predicted PDE terms. The inferred hidden physics term is given by,
@@
(a) Sample 1
(b) Sample 2
Figure 3: Reaction–diffusion equation: reference vs. DHPO predictions and HPM-integrated trajectories for two representative test samples drawn from the sine-function input space. Training data spans $t\leq 1$, while the extrapolation region ($t>1$) is indicated by white horizontal lines. For Sample 1, mean absolute errors are $u_{pred}(t\leq 1)=0.00432$, $u_{pred}(t>1)=0.04360$, $u_{hpm}(t\leq 1)=0.00226$, $u_{hpm}(t>1)=0.01917$. For Sample 2, mean absolute errors are $u_{pred}(t\leq 1)=0.01299$, $u_{pred}(t>1)=0.12473$, $u_{hpm}(t\leq 1)=0.00655$, $u_{hpm}(t>1)=0.05840$.
@@
Additionally, we assess the model’s performance on modified GRFs by varying the length-scale parameter $l$ of the RBF kernel with values $l\in\{0.1,0.15,0.20,0.40,0.60\}$. As expected, Table 2 demonstrates that prediction error increases as the length scale decreases. However, for out-of-distribution functions with boundary values approaching zero, the model maintains relatively higher accuracy compared to cases with non-zero boundary conditions.
  • Noise robustness:
    Considering the impact of noise on model performance, an assessment was conducted using training data corrupted with synthetically generated random uniform noise. The noise was introduced as follows $u_{noise}=u_{clean}+U(0,\gamma\sigma_{u})$ , where $U$ is sampled from uniform random distribution with minimum value 0, maximum $\gamma\sigma_{u}$ , $\sigma_{u}$ denotes standard deviation in clean data, the parameter $\gamma$ controls the noise intensity. We train the model considering $\gamma\in\{0,0.2,0.5\}$, for direct comparison, both the training data volume and the training methodology were kept identical to those utilized for the optimal model presented earlier. The results are computed considering 1000 test inputs from sine basis function space. As illustrated in Table 3, the test error predictably increases with rising noise intensity. Notably, for noise intensity $\gamma<0.2$, the error exhibits only a marginal increase, $\approx 10\%$ . This indicates that the model maintains robustness in capturing the underlying dynamics within this noise threshold.
  • Sparsity:
    To evaluate the model’s performance under conditions of data sparsity, training was conducted using only two extreme temporal data points (t=0and t=1)(t=0\ $\text${and }t=1) along with 101 spatial data points. We train the model with 500 samples using the similar training procedure as shown in the previously trained optimal model. Evaluating the model on 1000 test samples yielded a mean test error of $0.00681\pm 0.00337$. When compared against the optimal model which was trained on 500 spatio-temporal data points, these results demonstrate that the sparse data model is remarkably capable of accurately predicting the dynamics, despite its limited training data.
    To assess the benefits of physics-informed training, we compare against a standard DeepONet trained on the same sparse data using identical branch and trunk network architectures. Evaluating on 1000 test samples, the DeepONet achieves a prediction error of $0.02619\pm 0.01074$ at the training time instants but degrades to $0.43818\pm 0.21707$ at interior (unobserved) time instants. This indicates severe overfitting to the training timestamps with poor generalization to the interior time domain. Representative predictions are shown in Figure 10.
Figure 4: Performance analysis for Reaction diffusion equation: (a) Mean test error over varying $N_{\text{train}}$. (b) Test error distribution of optimal model that achieved an error $0.00871\pm 0.00424$ with $N_{train}=500$, $N_{d}=500$.
@@
0.10 0.07945 $\pm$ 0.03666
0.15 0.03496 $\pm$ 0.01831
0.20 0.01653 $\pm$ 0.00856
0.40 0.00755 $\pm$ 0.00421
0.60 0.00656 $\pm$ 0.00378
Table 3: Reaction Diffusion system: Mean and standard deviation of the test error over the noise level.
noise intensity $\gamma$ Test Error
0.0 0.00871 $\pm$ 0.00424
0.2 0.00983 $\pm$ 0.00515
0.5 0.01652 $\pm$ 0.00895
Figure 5: Performance analysis for GRFs and Modified GRFs input function spaces: (a) GRFs basis $l=0.2$, Test error = 0.03918 $\pm$ 0.02095. (b) Modified GRFs basis $l=0.2$, Test error = 0.01653 $\pm$ 0.00856
@@
Deterministic. The neural operator framework introduced in Section 2.2 is employed to characterize the system, specifically to identify the diffusion coefficient of the governing PDE using solution values at selected spatial and temporal locations.
@@
(a) Sample 1: Relative $L_{2}$ error of $u(x,t)$ = $0.027$. For this case, $D_{\text{true}}=0.027$ while $D_{\text{predicted}}=0.027$.
(b) Sample 2: Relative $L_{2}$ error of $u(x,t)$ = $0.029$. For this case, $D_{\text{true}}=0.036$ while $D_{\text{predicted}}=0.034$.
@@
Figure 7: Reaction Diffusion Equation: Distribution of absolute error of $D$ when compared to the $D$ used to generate the ground truth using the finite difference solver. The framework approximates $D$ with an error of $8\times 10^{-3}\pm 0.008$.
Probabilistic. To validate the probabilistic framework’s ability to capture parameter non-uniqueness, we evaluate whether parameter values sampled from the predicted distribution $\mathcal{N}(\mu_{D},\sigma_{D}^{2})$ can accurately reproduce the original sensor observations used as input to the inverse operator. For each test case, we sample 500 diffusion coefficient values from the learned distribution and perform forward simulations using a finite difference solver to obtain the corresponding solution fields. Figure 8 illustrates the predicted parameter distributions for two representative test samples, showing the mean estimate (dashed red line) alongside the true ground truth value (dashed green line). The distribution widths reflect the epistemic uncertainty inherent in the inverse problem given the available sparse sensor data. Figure 9 presents a scatter comparison between the true solution field $u$ (used to generate sensor measurements) and solutions obtained from sampled parameter values, demonstrating strong agreement across the spatial domain. Quantitatively, we achieve an average coefficient of determination $R^{2}=0.98$ across all sampled parameters, confirming that the inverse neural operator successfully captures the distribution of diffusion coefficients consistent with the observed data. This high reconstruction fidelity validates that the probabilistic framework appropriately quantifies parameter uncertainty while ensuring physical consistency.
Figure 8: Reaction Diffusion Equation: Predicted probability distributions of Diffusion coefficient $D$ for two representative test samples. The histogram represents 500 sampled values from the learned distribution $\mathcal{N}(\mu_{D},\sigma_{D}^{2})$. Dashed red line indicates the predicted mean, dashed green line shows the true Diffusion coefficient.
Figure 9: Reaction Diffusion Equation: Validation of predicted parameter distribution for representative samples through forward finite difference simulations. Each subplot shows the true solution field $u(x,t)$ (x-axis) versus the solution obtained using a Diffusion coefficient value sampled from the predicted distribution (y-axis).
3.2 Burgers’ equation
@@
[…]
@@
where the $\mathcal{N}$ denotes the hidden physics that is a function of the solution field $u(x,t)$ and its spatial derivatives $u_{x},u_{xx}$. In this problem, our aim is to learn the solution operator, $\mathcal{G}_{\theta}$ that maps the initial condition, $f(x,0)$ to the solution field, $u(x,t)$ using the partial known physics as well as sparse labeled dataset. In this case, we have considered fixed viscosity for the system, $\nu=0.01$. We solved equation (14) using a finite difference solver on a uniform grid ($\delta x=\delta t=0.01$), resulting in a $101\times 101$ grid spanning the $x$-$t$ domain with initial conditions sampled from a Gaussian Random Field (GRF).
The experiments were conducted with varying numbers of training samples, $N_{train}\in\{50,100,200,500,1000\}$ and the size of labeled dataset, $N_{d}\in\{200,500\}$, where $N_{d}$ denotes the number of locations per sample where the value of solution field is known. For testing, we used a fixed set of $N_{test}=1,000$ samples. Figure 10(a) illustrates the mean test error across different combinations of sample sizes and training data points. The optimal performance was achieved with $N_{train}=1000$ and $N_{d}=500$, resulting in mean absolute error of $0.00680\pm 0.00564$. The loss curve over the training process is shown in Figure 1(b).
Figure 10: Performance analysis for Burger’s equation: (a) Mean test error over varying $N_{\text{train}}$. (b) Test error distribution of optimal model that achieved an error $0.00680\pm 0.00564$ with $N_{train}=1000$, $N_{d}=500$.
The distribution of test errors for this optimal configuration is presented in Figure 10(b). As demonstrated in Figure 11, the model accurately predicts the solution when compared against ground truth. We also observe, in extrapolation region, the model accurately predicts the solution up to a specific time threshold, Hence we can infer that the Hidden Physics MLP is accurately capturing the dynamics, The Figure 9 shows the contours of PDE reference and predicted PDE terms, the model successfully approximated the unknown hidden term as,
@@
(a) Sample 1
(b) Sample 2
Figure 11: Burgers’ equation: reference vs. DHPO predictions and HPM-integrated trajectories for two representative test samples. Training data spans $t\leq 1$, and the extrapolation region ($t>1$) is indicated by white horizontal lines. For Sample 1, mean absolute errors are $u_{pred}(t\leq 1)=0.00250$, $u_{pred}(t>1)=0.00883$, $u_{hpm}(t\leq 1)=0.00456$, $u_{hpm}(t>1)=0.01198$. For Sample 2, mean absolute errors are $u_{pred}(t\leq 1)=0.01268$, $u_{pred}(t>1)=0.02325$, $u_{hpm}(t\leq 1)=0.01454$, $u_{hpm}(t>1)=0.02128$.
Relation to symbolic regression. For comparison, we applied weak-SINDy Messenger and Bortz [2021], a sparse symbolic regression method, to 300 randomly sampled measurements from a single trajectory, recovering $u_{t}=0.0101\cdot u_{xx}-1.0068\cdot u\cdot u_{x}$ with residual $3.95\times 10^{-3}$. While weak-SINDy efficiently discovers symbolic equations from limited data, DHPO learns generalizable operators of solution across varying initial conditions with physics discovery as an integrated component. The methods address complementary aspects of the inverse problem.
@@
Deterministic. We are now interested in employing the neural operator framework discussed in Section 2.2 to characterize the system, more specifically to identify the viscosity of the model while being provided with the solution values at a few spatial and temporal locations.